# Cycle Matroids of Graphings: From Convergence to Duality

### v2

@@ 4.2 Weakly exposed points
We say that an extreme point $x\in C$ is exposed in two steps by $L^{1}$ functions, if one can find $g_{1},g_{2}\in L^{1}(E,\tilde{\mu})$ such that $\varphi_{g_{1}}(x)$ is maximal in $C$, and $x$ to strictly maximizes $\varphi_{g_{2}}$ in $C^{\prime}=\{y\in C\mid\varphi_{g_{1}}(y)=\varphi_{g_{1}}(x)\}$. Note that the proof of Theorem 4.4 yields the same conclusion for points that are exposed in two steps. The only difference is that during the algorithm, we always choose an edge that is maximal with respect to $g_{1}$, and in case of ties, we choose the one maximal with respect to $g_{2}$. Finally, if there is still a tie, we decide according to the Borel ordering we fixed at the beginning.
More generally, let $\alpha\in\omega_{1}$ be a countable ordinal and let $\mathbf{g}\colon\alpha\to L^{1}(E,\tilde{\mu})$ be a list of functions. Let us define $\mathbf{g}(x)\in\mathbb{R}^{\alpha}$ by $f(x)_{\beta}=\mathbf{g}(\beta)(x)$. Then, $\mathbf{g}$ defines a Borel partial ordering on $C$ based on the lexicographical ordering of the images of points of $C$ in $\mathbb{R}^{\alpha}$. We denote this ordering by $<_{\mathbf{g}}$ and call a point $x\in C$ exposed in $\alpha$ steps by the list $\mathbf{g}$ if $x$ is the unique maxima of $<_{\mathbf{g}}$ in $C$. In other words, for all $y\in C$ there exists an ordinal $\beta_{y}\leq\alpha$ such that for all $\gamma\leq\beta_{y}$, we have $\mathbf{g}(\gamma)(x)=\mathbf{g}(\gamma)(y)$ and $\mathbf{g}(\beta_{y})(x)>\mathbf{g}(\beta_{y})(y)$. We say $x\in C$ is weakly exposed, if it is exposed in $\alpha$ steps for some countable ordinal $\alpha$.
@@
Let $f_{\alpha}\in C$ be weakly exposed by a list of functions in $L^{1}(E,\tilde{\mu})$. Then, $f_{\alpha}$ is 00-$1$ valued $\tilde{\mu}$-almost everywhere.
@@
Weakly exposed points of $C$ are strongly exposed.
In finite dimension, all extreme points of a compact convex set are weakly exposed. In infinite dimension, however, this is not necessarily true. The closed unit ball of $\ell^{p}(\mathbb{N})$ for $21<p<$ 2 in $\ell^{2}(\mathbb{N})$ has extreme points that are not even support points, see [23, Example 7.5].
@@
The algorithm in the proof of Theorem 4.4 uses an arbitrary Borel ordering $<$ on the edge set. The resulting forest is called the Wired maximal spanning forest, see  [20, Chapter 11]. Usually, such forests are studied with the ordering coming from iid $[0,1]$ labels on the edges of a countable graph.
@@ 5 Duality
Roughly speaking, a planar graphing is a graphing $\mathcal{G}=(X,E,\mu)$ whose connected components are planar and are also endowed with a concrete embedding in a measurable way. The embedding is encoded by specifying the finite facial cycles of the components, and we assume this collection of edge sets to be Borel inside the standard Borel space of finite subsets of $E$. For the precise definition, see [7, Section 3.2]. The component of a $\mu$-random vertex, together with the embedding is a unimodular random map in the sense of [1].
We will, in fact, restrict our attention to the case when all facial cycles of $\mathcal{G}$ are finite, and refer to them as faces. The assumption of $2$-vertex-connectivity and bounded size of faces mentioned in the Introduction also applies. The fact that the encoding is Borel ensures that we can talk about the dual graphing $\mathcal{G}^{*}=(X^{*},E^{*},\mu^{*})$, where $X^{*}$ is the set of faces of $\mathcal{G}$. There is a measure preserving Borel bijection $\sigma\colon(E,\tilde{\mu})\to(E^{*},\tilde{mu}^{*})$ that denotes the usual correspondence: $\sigma(e)$ connects two distinct faces $x_{1}^{*}$ and $x_{2}^{*}$ if and only if $e$ is a common boundary edge of $x_{1}^{*}$ and $x_{2}^{*}$. For sake of simplicity, we write $e^{*}=\sigma(e)$ and $F^{*}=\sigma(F)$ for $F\subseteq E(\mathcal{G})$. For an edge set $F\subseteq E$, we use the notation $F^{*}=\{e^{*}\mid e\in F\}$.
Examples of planar graphings include Schreier graphings of p.m.p. actions of finitely generated groups with planar Cayley graphs (in particular surface groups), as well as Palm graphings of planar factor graphs of free invariant point processes on $\mathbb{R}^{2}$ or $\mathbb{H}^{2}$ [21, Remark 17].
@@
For graphings, we usually assume $\mu(X)=1$ for convenience. For finite graphs this corresponds to normalization by the number of vertices. However, when duals are considered, this normalization is less natural, since the dual graph typically does not have the same number of vertices. It would probably be more natural to normalize by the number of edges which is the same for the two graphs. Nevertheless, to be consistent with the rest of the paper, we do not change the normalization, and simply accept that the vertex measure of the dual graphing is not necessarily 1.
@@
We define the cocyclic matroid of $\mathcal{G}$ by the infinite analogue of the usual rank function formula:
$\rho^{D}_{\mathcal{G}}(A)=\rho_{\mathcal{G}}(E\setminus A)+\tilde{\mu}(A)-\rho_{\mathcal{G}}(E).$
Note that $\rho_{\mathcal{G}}(E)=1$, and $\rho^{D}_{\mathcal{G}}$ is submodular as it is the sum of a submodular and a modular function. Moreover, the rank of the full edge set in the cocycle matroid is $\rho^{D}_{\mathcal{G}}(E)=\tilde{\mu}(E)-\mu(X)$. There is a nice correspondence between the elements of the associated base-polytopes:
$\alpha\in\text{bmm}^{+}(\rho_{\mathcal{G}})\iff\tilde{\mu}-\alpha\in\text{bmm}^{+}(\rho^{D}_{\mathcal{G}}).$
@@ 5.1 Duality for hyperfinite planar graphings
For finite planar graphs, the cocycle matroid is the cycle matroid of the dual. In the Borel setting this holds only for hyperfinite planar graphings. In fact, we will exploit exactly the discrepancy between $\rho_{\mathcal{G}}$ and $\rho^{D}_{\mathcal{G}^{*}}$ in Subsection 5.2 to prove Theorem 1.9.
@@
Theorem 1.8 is actually a reformulation of Euler’s formula, stating that for any embedding of a finite planar graph, the equality $n-m+f=2$ holds, where $n$, $m$ and $f$ denote the number of vertices, edges and faces, respectively. For an planar graphing $\mathcal{G}$ with infinite components, the special case $\rho^{D}_{\mathcal{G}*}(E)=\rho_{\mathcal{G}}(E)$ translates into $\tilde{\mu}(E)-\mu^{*}(X^{*})=\mu(X)$, that is, the $2$ in the formula disappears in the limit.
@@
We noted earlier that $\text{bmm}^{+}(\rho^{D}_{\mathcal{G}^{*}})=\tilde{\mu}-\text{bmm}^{+}(\rho_{\mathcal{G}^{*}})$. By Theorem 4.4, we know that exposed points of $\text{bmm}^{+}(\rho_{\mathcal{G}})$ and $\text{bmm}^{+}(\rho_{\mathcal{G}^{*}})$ are essential hyperfinite spanning forests of $\mathcal{G}$ and $\mathcal{G}^{*}$, respectively. Note also that they are in a one-to-one correspondence by complementation. Here we use that $\mathcal{G}$ and $\mathcal{G}^{*}$ are hyperfinite, to ensure that the complement is again hyperfinite.
These observations together show that $\text{bmm}^{+}(\rho_{\mathcal{G}})$ and $\text{bmm}^{+}(\rho^{D}_{\mathcal{G}^{*}})$ have the same set of exposed points. This implies $\text{bmm}^{+}(\rho_{\mathcal{G}})=\text{bmm}^{+}(\rho^{D}_{\mathcal{G}^{*}})$, as they are the weak* closure of the convex hull of the same set, see Remark 4.7. Finally, this implies $\rho_{\mathcal{G}}=\rho^{D}_{\mathcal{G}^{*}}$ by the fact that a submodular function $\varphi$ is determined by $\text{bmm}^{+}(\varphi)$ in the sense that $\varphi(A)=\sup_{\alpha\in\text{bmm}^{+}(\varphi)}\alpha(A)$, see [19, Corollary 6.10]. The proof of $\rho^{D}_{\mathcal{G}}=\rho_{\mathcal{G}^{*}}$ is analogous. ∎
@@
By using the definitions of $\rho_{\mathcal{G}}$ and $\rho^{D}_{\mathcal{G}^{*}}$, and the equalities $\tilde{\mu}^{*}=\tilde{\mu}$ and $\rho_{\mathcal{G}^{*}}(E^{*})=\mu^{*}(X^{*})$, we obtain that $\rho_{\mathcal{G}}(A)=\rho_{\mathcal{G}^{*}}(A)$ is equivalent to
[…]
To verify the equality, we set up a payment between vertices, faces, and edges. Motivated by the second integral, every face $x^{*}$ that is in a finite $(E^{*}\setminus A^{*})$-component pays $1/|V(\mathcal{G}^{*,E^{*}\setminus A^{*}}_{x^{*}})|$ token to the component. This way every finite $(E^{*}\setminus A^{*})$-component collects 1 token in total, and this token is passed on to an edge (chosen in a Borel way) in the $A$-cycle surrounding the component. We can also make sure that the edges that get payed, oriented from the component that pays to them, form an acyclic oriented graph on the components of $E^{*}\setminus A^{*}$. This way the edges of $A$ that do not receive a payment form a subforest $A^{\prime}\subseteq A$ that has the same connected components in $\mathcal{G}$ as $A$.
Next, every $x\in X$ in a finite $A$-component pays $1/|V(\mathcal{G}^{A}_{x})|$ token to every $A^{\prime}$-edge in its component. This way all such vertices pay $1-1/|V(\mathcal{G}^{A}_{x})|$, and such $A$-edges receive $1$ token in total.
What remains are infinite $A^{\prime}$ components and their vertices. Write $A^{\prime\prime}$ for this edge set, and $X^{\prime\prime}$ for the vertices. Then $\tilde{\mu}(A^{\prime\prime})=\mu(X^{\prime\prime})$ by Proposition 2.2, as $A^{\prime\prime}$ is a hyperfinite treeing on $X^{\prime\prime}$ with infinite components. In a mass transport formulation, this means that we can arrange that every vertex in $X^{\prime\prime}$ pays 1 token, and every edge in $A^{\prime\prime}$ gets 1 token.
Concluding the above, vertices and faces payed in total the right hand side of (4), while every edge in $A$ received 1, so (4) holds by the Mass Transport Principle. ∎

### v3

@@ 4.3 Weakly exposed points
We say that an extreme point $x\in\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ is exposed in two steps by $L^{1}$ functions, if one can find $g_{1},g_{2}\in L^{1}(E,\eta_{\mathbf{G}})$ such that $\varphi_{g_{1}}(x)$ is maximal in $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$, and $x$ strictly maximizes $\varphi_{g_{2}}$ in $\{y\in\text{\sf bmm}_{+}(\rho_{\mathbf{G}})\mid\varphi_{g_{1}}(y)=\varphi_{g_{1}}(x)\}$. Note that the proof of Theorem 4.2 yields the same conclusion for points that are exposed in two steps. The only difference is that during the algorithm, we always choose an edge that is maximal with respect to $g_{1}$, and in case of ties, we choose the one maximal with respect to $g_{2}$. Finally, if there is still a tie, we decide according to the Borel ordering we fixed at the beginning.
More generally, let $\alpha\in\omega_{1}$ be a countable ordinal and let $\mathbf{g}\colon\alpha\to L^{1}(E,\eta_{\mathbf{G}})$ be a list of functions. Let us define $\mathbf{g}(x)\in\mathbb{R}^{\alpha}$ by $f(x)_{\beta}=\mathbf{g}(\beta)(x)$. Then, $\mathbf{g}$ defines a Borel partial ordering on $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ based on the lexicographical ordering of the images of points of $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ in $\mathbb{R}^{\alpha}$. We denote this ordering by $<_{\mathbf{g}}$ and call a point $x\in\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ exposed in $\alpha$ steps by the list $\mathbf{g}$ if $x$ is the unique maxima of $<_{\mathbf{g}}$ in $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$. In other words, for all $y\in\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ there exists an ordinal $\beta_{y}\leq\alpha$ such that for all $\gamma\leq\beta_{y}$, we have $\mathbf{g}(\gamma)(x)=\mathbf{g}(\gamma)(y)$ and $\mathbf{g}(\beta_{y})(x)>\mathbf{g}(\beta_{y})(y)$. We say $x\in\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ is weakly exposed, if it is exposed in $\alpha$ steps for some countable ordinal $\alpha$.
@@
Let $f_{\alpha}\in\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ be weakly exposed by a list of functions in $L^{1}(E,\eta_{\mathbf{G}})$. Then, $f_{\alpha}$ is 00-$1$ valued $\eta_{\mathbf{G}}$-almost everywhere.
@@
Weakly exposed points of $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ are strongly exposed.
In finite dimension, all extreme points of a compact convex set are weakly exposed. In infinite dimension, however, this is not necessarily true. The closed unit ball of $\ell^{p}(\mathbb{N})$ for $21<p<$ 2 in $\ell^{2}(\mathbb{N})$ has extreme points that are not even support points, see [28, Example 7.5].
@@
The algorithm in the proof of Theorem 4.2 results in a Wired Maximal Spanning Forest with respect to an arbitrary Borel ordering $<$ on the edge set, as defined in the introduction. Usually, such forests are studied with the ordering coming from iid $[0,1]$ labels on the edges of a countable graph.
4.4 Characterizations of independent sets and bases
In this subsection we prove Theorems 1.2 and 1.3.
Proof of Theorem 1.2.
First we prove (a). For the ‘if’ direction, note that $\eta_{\mathbf{G}}|_{F}\in\text{\sf mm}_{+}(\rho_{\mathbf{G}})$ by Lemma 4.1. The Radon-Nikodym derivative of the measure $\eta_{\mathbf{G}}|_{F}$ is ${\mathbbm{1}}_{F}$. The function $g={\mathbbm{1}}_{F}-{\mathbbm{1}}_{F^{c}}$ is a witness function for $\alpha$ being exposed. To see the ‘only if’ direction, using Theorem 4.4, we know that exposed functions are $0-1$ valued. Therefore, the exposed points are measures of the form $\eta_{\mathbf{G}}|_{F}$ for some subset $F\subseteq E$. However, the subsets for which this restriction is in $\text{\sf mm}_{+}$ are exactly the hyperfinite subforests by Lemma 4.1.
Now we prove (b). For the ‘if’ direction, note that $\eta_{\mathbf{G}}|_{F}\in\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ by Lemma 4.1. The Radon-Nikodym derivative of the measure $\eta_{\mathbf{G}}|_{F}$ is ${\mathbbm{1}}_{F}$. The function $g={\mathbbm{1}}_{F}-{\mathbbm{1}}_{F^{c}}$ is a witness function for $\alpha$ being exposed. To see the ‘only if’ direction, using Theorem 4.4, we know the exposed functions are $0-1$ valued. Therefore, the exposed points are measure of the form $\eta_{\mathbf{G}}|_{F}$ for some subset $F\subseteq E$. However, the subsets for which this restriction is in $\text{\sf bmm}_{+}(\rho_{\mathbf{G}})$ are exactly the hyperfinite essential spanning forests by Lemma 4.1. ∎
Proof of Theorem 1.3.
[…]
4.5 Independent sets and exchange properties
The setfunction $\rho_{\mathbf{G}}$ was shown to be submodular in [23]. The proof relies on structural analysis of so-called $\mathcal{B}$-partitions and is somewhat complicated. In this subsection we give a new proof of this result that conceptually differs from the original one. The main idea is to define a submodular setfunction through independent sets, and then to show that the function thus obtained is identical to $\rho_{\mathbf{G}}$.
Let $\mathbf{G}=(J,\mu,E)$ be a graphing. We define $\mathcal{I}(\mathbf{G})=\{F\subseteq E\mid F\text{ is a hyperfinite forest}\}$ and $\mathcal{B}(\mathbf{G})=\{F\subseteq E\mid F\text{ is a hyperfinite essential spanning forest}\}$. By Corollary 4.3, $\mathcal{I}(\mathbf{G})$ is the set of subsets of the elements of $\mathcal{B}(\mathbf{G})$, in other words,
$\mathcal{I}(\mathbf{G})=\{F\mid F\subseteq F^{\prime}\ \text{for some}\ F^{\prime}\in\mathcal{B}(\mathbf{G})\}.$
By Lemma 2.1, the measure of any hyperfinite essential spanning subforest is the same.
Let $(J,\mu)$ be a standard Borel space. We say that family of subsets $\mathcal{I}$ of $J$ is the independence family of a measurable matroid with respect to the measure $\mu$, if
  1. (I0)
    $\mathcal{I}$ is closed under taking subsets,
  2. (I1)
    if $I_{1}\subseteq I_{2}\subseteq\dots$ with $I_{n}\in\mathcal{I}$ for all $n\in\mathbb{Z}_{+}$, then $\cup I_{n}\in\mathcal{I}$,
  3. (I2)
    for all $I_{1},I_{2}\in\mathcal{I}$ with $\mu(I_{1})<\mu(I_{2})$, there exists $I_{3}\in\mathcal{I}$ such that $I_{1}\subseteq I_{3}\subseteq I_{1}\cup I_{2}$ and $\mu(I_{3})>\mu(I_{1})$,
In certain cases, the following reformulation of (I2) is easier to work with.
  1. (I2’)
    for all $I_{1},I_{2}\in\mathcal{I}$, there exists $I_{3}\in\mathcal{I}$ such that $I_{1}\subseteq I_{3}\subseteq I_{1}\cup I_{2}$ and $\mu(I_{3})\geq\mu(I_{2})$.
Lemma 4.10.
[…]
Proof.
It is clear that (I2’) implies (I2). It remains to show that (I1) and (I2) together imply (I2’). We build a transfinite sequence of Borel sets $A_{\alpha}$ ($\alpha\leq\omega_{1}$). Let $A_{0}=I_{1}$. Using (I2), there exists $A_{1}\in\mathcal{I}$ such that $A_{0}\subseteq A_{1}\subseteq A_{0}\cup I_{2}$. Similarly, there exists $A_{2}\in\mathcal{I}$ such that $A_{1}\subseteq A_{2}\subseteq A_{1}\cup I_{2}=A_{0}\cup I_{2}$. In general, let $f\colon\mathcal{I}\to\mathcal{I}$ be a function with the following properties: $A\subseteq f(A)\subseteq I_{2}\cup A$ for all $A\in\mathcal{I}$, $\mu(A)<\mu(f(A))$ whenever $\mu(A)<\mu(I_{2})$, and $A=f(A)$ whenever $\mu(A)=\mu(I_{2})$. Note that such a function exists by (I2). Now do this recursion transfinitely, up to ordinal $\omega_{1}$.
If $\mu(A_{\alpha})=\mu(I_{2})$ for some $A_{\alpha}$ in the sequence, then we are done. Suppose that $\mu(A_{\alpha})<\mu(I_{2})$ for all $A_{\alpha}$. Since there exists no increasing sequence of real numbers of order $\omega_{1}$, the sequence $A_{\alpha}$ has to stabilise at a certain point. Call the first ordinal with such an independent set $\alpha_{0}$; note that $A_{\alpha_{0}}$ is in $\mathcal{I}$ by (I1). That is, $f(A_{\alpha_{0}})=A_{\alpha_{0}}$, but $\mu(A_{\alpha_{0}})<\mu(I_{2})$, contradicting the choice of $f$.
Lemma 4.11.
Let $\mathcal{I}$ be the family of independent subsets of a measurable matroid with respect to measure $\mu$. Then the setfunction $r(A)=\sup\{\mu(I)\mid I\in\mathcal{I},\ I\subseteq A\}$ is submodular.
Proof.
We here imitate the proof of [22, Theorem 5.12]. Let $A,B$ Borel sets and $\varepsilon>0$. Then there exists $I_{1},I_{2}\in\mathcal{I}$ such that $I_{1}\subseteq A\cap B,I_{2}\subseteq A\cup B$, and $r(A\cap B)-\varepsilon\leq\mu(I_{1})$, $r(A\cup B)-\varepsilon\leq\mu(I_{2})$. By the exchange property (I2’), there exists a set $H\in\mathcal{I}$ such that $I_{1}\subseteq H\subseteq I_{1}\cup I_{2}\subseteq A\cup B$ and $\mu(H)\geq\mu(I_{2})\geq\mu(A\cup B)-\varepsilon$. Then $\mu(H\cap(A\cap B))\geq\mu(I_{1}\cap(A\cap B))=\mu(I_{1})\geq r(A\cap B)-\varepsilon$. Therefore, using (I0), we get
$r(A)+r(B){}$ ${}\geq r(H\cap A)+r(H\cap B)$
${}=\mu(H\cap A)+\mu(H\cap B)$
${}=\mu(H\cap(A\cap B))+\mu(H\cap(A\cup B))$
${}\geq r(A\cap B)-\varepsilon+\mu(A\cup B)-\varepsilon.$
As this inequality holds for all $\varepsilon>0$, it follows that $r$ is submodular. ∎
With the help of Lemma 4.11, first we show that $\mathcal{I}(\mathbf{G})$ defines a submodular function.
Lemma 4.12.
Let $\mathbf{G}$ be a graphing. Then $\mathcal{I}(\mathbf{G})$ forms the family of independent sets of a measurable matroid with respect to the edge measure $\eta_{\mathbf{G}}$.
Proof.
It is not difficult to check that Properties (I0) and (I1) are satisfied, hence we concentrate on (I2’). Let $I_{1},I_{2}\in\mathcal{I}(Gb)$ and consider the graphing $\mathbf{H}=(V,\mu,I_{1}\cup I_{2})$. Then $\eta_{\mathbf{H}}=\eta_{\mathbf{G}}|_{I_{1}\cup I_{2}}$. Let $I_{3}$ be an extension of $I_{1}$ such that $I_{3}$ is a hyperfinite essential spanning forest in $\mathbf{H}$; such an extensions exists by Corollary 4.3. Then $\eta_{\mathbf{G}}(I_{3})=\eta_{\mathbf{G}}(I_{1}\cup I_{2})\geq\eta_{\mathbf{G}}(I_{2})$, therefore the exchange property (I2’) holds. ∎
Combining the above results, now we are ready to prove that $\rho_{\mathbf{G}}$ is submodular.
Theorem 4.13.
Let $\mathbf{G}=(J,\mu,E)$ be a graphing. Then the setfunction $\rho_{\mathbf{G}}$ defined by (1) is submodular.
Proof.
Every graphing contains a hyperfintie essential spanning forest. Furthermore, by Corollary 2.2, the edge measure of a hyperfinite essential spanning forest is exactly the rank of the graphing. From these observations it follows that for any $A\subseteq E$, $\rho_{\mathbf{G}}(A)=\sup\{\eta_{\mathbf{G}}(F)\mid F\subseteq A\ \text{is a hyperfinite forest in}\ (J,\mu,A)\}$. By Lemma 4.12, hyperfinite subforests form the family of independents sets of a measurable matroid with respect to the edge measure $\eta_{\mathbf{G}}$. Therefore, the theorem follows by Lemma 4.11. ∎
@@ 5 Duality
Roughly speaking, a planar graphing is a graphing $\mathbf{G}=(J,\mu,E)$ whose connected components are planar and are also endowed with a concrete embedding in a measurable way. The embedding is encoded by specifying the finite facial cycles of the components, and we assume this collection of edge sets to be a Borel set inside the standard Borel space of finite subsets of $E$. For the precise definition, see [10, Section 3.2]. The component of a $\mu$-random vertex, together with the embedding is a unimodular random map in the sense of [3].
We will, in fact, restrict our attention to the case when all facial cycles of $\mathbf{G}$ are finite, and refer to them as faces. The assumption of $2$-vertex-connectivity and bounded size of faces mentioned in the Introduction also applies. The fact that the encoding is Borel ensures that we can talk about the dual graphing $\mathbf{G}^{*}=(J^{*},\mu^{*},E^{*})$, where $J^{*}$ is the set of faces of $\mathbf{G}$. There is a measure preserving Borel bijection $\sigma\colon E\to E^{*}$ that denotes the usual correspondence: $\sigma(e)$ connects two distinct faces $x_{1}^{*}$ and $x_{2}^{*}$ if and only if $e$ is a common boundary edge of $x_{1}^{*}$ and $x_{2}^{*}$. For sake of simplicity, we write $e^{*}=\sigma(e)$. For an edge set $F\subseteq E$, we use the notation $F^{*}=\{e^{*}\mid e\in F\}$.
Examples of planar graphings include Schreier graphings of p.m.p. actions of finitely generated groups with planar Cayley graphs (in particular surface groups), as well as Palm graphings of planar factor graphs of free invariant point processes on $\mathbb{R}^{2}$ or $\mathbb{H}^{2}$ [25, Remark 17].
@@
For graphings, we usually assume $\mu(J)=1$ for convenience. For finite graphs this corresponds to normalization by the number of vertices. However, when duals are considered, this normalization is less natural, since the dual graph typically does not have the same number of vertices. It would probably be more natural to normalize by the number of edges which is the same for the two graphs. Nevertheless, to be consistent with the rest of the paper, we do not change the normalization, and simply accept that the vertex measure of the dual graphing is not necessarily 1.
@@
We define the cocyclic matroid of $\mathbf{G}$ by the infinite analogue of the usual dual rank function formula:
$\rho^{*}_{\mathbf{G}}(F)=\rho_{\mathbf{G}}(E\setminus F)+\eta_{\mathbf{G}}(F)-\rho_{\mathbf{G}}(E).$
Note that $\rho^{*}_{\mathbf{G}}$ is submodular as it is the sum of a submodular and a modular function. Furthermore, $\rho_{\mathbf{G}}(E)\leq 1$, and equality holds if and only if almost all components of $\mathbf{G}$ are infinite. Moreover, the rank of the full edge set in the cocycle matroid is $\rho^{*}_{\mathbf{G}}(E)=\eta_{\mathbf{G}}(E)-1$. There is a nice correspondence between the elements of the associated minorizing measures:
$\alpha\in\text{\sf bmm}^{+}(\rho_{\mathbf{G}})\iff\eta_{\mathbf{G}}-\alpha\in\text{\sf bmm}^{+}(\rho^{*}_{\mathbf{G}}).$
@@ 5.1 Duality for hyperfinite planar graphings
For finite planar graphs, the cocycle matroid is the cycle matroid of the dual. In the Borel setting, this holds only for hyperfinite planar graphings. In fact, we will exploit exactly this discrepancy between $\rho_{\mathbf{G}}$ and $\rho^{*}_{\mathbf{G}^{*}}$ in Subsection 5.2 to prove Theorem 1.7.
@@
Theorem 1.6 is actually a reformulation of Euler’s formula, stating that for any embedding of a finite planar graph, the equality $n+f=m+2$ holds, where $n$, $m$ and $f$ denote the number of vertices, edges and faces, respectively. For a planar graphing $\mathbf{G}$ with infinite components, the special case $\rho^{*}_{\mathbf{G}^{*}}(E^{*})=\rho_{\mathbf{G}}(E)$ translates into $\mu(J)+\mu^{*}(J^{*})=\eta_{\mathbf{G}}(E)$, that is, the $2$ in the formula disappears in the limit.
@@
First we prove $\rho^{*}_{\mathbf{G}^{*}}=\rho_{\mathbf{G}}$. We noted earlier that $\text{\sf bmm}^{+}(\rho^{*}_{\mathbf{G}^{*}})=\eta_{\mathbf{G}^{*}}-\text{\sf bmm}^{+}(\rho_{\mathbf{G}^{*}})$. By Theorem 4.2, we know that exposed points of $\text{\sf bmm}^{+}(\rho_{\mathbf{G}})$ and $\text{\sf bmm}^{+}(\rho_{\mathbf{G}^{*}})$ are hyperfinite essential spanning forests of $\mathbf{G}$ and $\mathbf{G}^{*}$, respectively. Note also that they are in a one-to-one correspondence by complementation. Here we use that $\mathbf{G}$ and $\mathbf{G}^{*}$ are hyperfinite, to ensure that the complement is again hyperfinite.
These observations together show that $\text{\sf bmm}^{+}(\rho_{\mathbf{G}})$ and $\text{\sf bmm}^{+}(\rho^{*}_{\mathbf{G}^{*}})$ have the same set of exposed points. This implies $\text{\sf bmm}^{+}(\rho_{\mathbf{G}})=\text{\sf bmm}^{+}(\rho^{*}_{\mathbf{G}^{*}})$, as they are the weak* closure of the convex hull of the same set, see Remark 4.6. Finally, this implies $\rho_{\mathbf{G}}=\rho^{*}_{\mathbf{G}^{*}}$ by the fact that a submodular function $\varphi$ is determined by $\text{\sf bmm}^{+}(\varphi)$ in the sense that $\varphi(A)=\sup_{\alpha\in\text{\sf bmm}^{+}(\varphi)}\alpha(A)$, see [22, Corollary 6.10]. The proof of $\rho^{*}_{\mathbf{G}}=\rho_{\mathbf{G}^{*}}$ is analogous. ∎
@@
By using the definitions of $\rho_{\mathbf{G}}$ and $\rho^{*}_{\mathbf{G}^{*}}$ and the equalities $\eta_{\mathbf{G}^{*}}=\eta_{\mathbf{G}}$ and $\rho_{\mathbf{G}^{*}}(E^{*})=\mu^{*}(J^{*})$, we obtain that $\rho_{\mathbf{G}}(F)=\rho^{*}_{\mathbf{G}^{*}}(F)$ is equivalent to
$\eta_{\mathbf{G}}(F)=\int_{x\in J}1-\frac{1}{|V(\mathbf{G}^{F}_{x})|}\mathop{}\!\mathrm{d}\mu(x)+\int_{x^{*}\in J^{*}}\frac{1}{|V(\mathbf{G}^{*,E^{*}\setminus F^{*}}_{x^{*}})|}\mathop{}\!\mathrm{d}\mu^{*}(x^{*}).$ (2)
To verify the equality, we set up a payment between vertices, faces, and edges. Motivated by the second integral, every face $x^{*}$ that is in a finite $(E^{*}\setminus F^{*})$-component pays $1/|V(\mathbf{G}^{*,E^{*}\setminus F^{*}}_{x^{*}})|$ token to the component. This way every finite $(E^{*}\setminus F^{*})$-component collects 1 token in total, and this token is passed on to an edge (chosen in a Borel way) in the $F$-cycle surrounding the component. We can also make sure that the edges that get payed, oriented from the component that pays to them, form an acyclic oriented graph on the components of $E^{*}\setminus F^{*}$. This way the edges of $F$ that do not receive a payment form a subforest $F^{\prime}\subseteq F$ that has the same connected components in $\mathbf{G}$ as $F$.
Next, every $x\in J$ in a finite $F$-component pays $1/|V(\mathbf{G}^{F}_{x})|$ token to every $F^{\prime}$-edge in its component. This way all such vertices pay $1-1/|V(\mathbf{G}^{F}_{x})|$, and such $F$-edges receive $1$ token in total.
What remains are infinite $F^{\prime}$ components and their vertices. Write $F^{\prime\prime}$ for this edge set, and $J^{\prime\prime}$ for the vertices. Then $\eta_{\mathbf{G}}(F^{\prime\prime})=\mu(J^{\prime\prime})$ by Theorem 2.1, as $F^{\prime\prime}$ is a hyperfinite treeing on $J^{\prime\prime}$ with infinite components. In a mass transport formulation, this means that we can ensure that every vertex in $J^{\prime\prime}$ pays 1 token, and every edge in $F^{\prime\prime}$ gets 1 token.
Concluding the above, vertices and faces payed in total the right hand side of (2), while every edge in $F$ received 1, so (2) holds by the Mass Transport Principle. ∎